OpenStudy (anonymous):

find the LCD of booth fractions and then simplify the expression completly assume the denominator equils 0 2x 3y ____ - ______ 6x^3y 4xy^2

OpenStudy (anonymous):

this could help with finding the LCD http://www.helpwithfractions.com/math-homework-helper/least-common-denominator/

OpenStudy (anonymous):

ok let me check it out

OpenStudy (anonymous):

can you do live help please

OpenStudy (anonymous):

live help?

OpenStudy (anonymous):

yea like help me now walk through it with me it is easyer

OpenStudy (anonymous):

okay. well for this question, when you say to assume that the denominator equals zero, does that mean to just completely avoid the bottom of each fractions?

OpenStudy (anonymous):

idk

OpenStudy (michele_laino):

we can rewrite your expression as below: \[\Large \frac{{2x}}{{6{x^3}y}} - \frac{{3y}}{{4x{y^2}}} = \frac{1}{{3{x^2}y}} - \frac{3}{{4xy}}\]

OpenStudy (anonymous):

what did u take out?

OpenStudy (anonymous):

The LCD is the smallest number or expression that is divisible by both denominators. You need to take each part of it at a time. Let's look at the numbers, then the x's then the y's. What is the smallest number that is divisible by both 3 and 6. In other words, what is the LCM of 3 and 6?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

No, it is 6

OpenStudy (michele_laino):

we have to compute the LCM between: \[\Large 3{x^2}y,\quad and\quad 4xy\]

OpenStudy (anonymous):

This means that the number part we need for the LCD is 6. Now let's look at the x's. There is an x^2 and an x. The lowest expression that will divide both x^2 and x is x^2, so we need x^2 for the LCD. Now let's look at the y part. You have y and y^3. The simplest expression with y that will divide evenly by y and y^3 is y^3, so you need y^3 in the LCD. Once you put together the number part, the x part and rthe y part, you end up with: LCD = 6x^2y^3

OpenStudy (anonymous):

I am so confused I have 2 diffrent ppl telling me 2 different things

OpenStudy (anonymous):

I'm not sure where the other person is going with their answer but mine are clear and simple

OpenStudy (anonymous):

ok

OpenStudy (michele_laino):

I'm sure that my procedure is the right procedure

OpenStudy (anonymous):

.... aren't you a qualified helper lol

OpenStudy (anonymous):

ook

OpenStudy (anonymous):

but mine is correct I'm sure

OpenStudy (michele_laino):

ok! Then please continue! @aaldia678

OpenStudy (anonymous):

I want to know what each of ur answers r

OpenStudy (anonymous):

@Michele_Laino thank you

OpenStudy (anonymous):

reread what I wrote @sarahefal

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wait I'm sorry, I used the wrong number

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

this is your equation right? \[\frac{ 2 }{ 6x ^{3y} } -\frac{ 3y }{ 4xy ^{2}}\]

OpenStudy (anonymous):

okay so what is the LCD of 6 and 4? figure that out first

OpenStudy (anonymous):

3*2^2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

the LCD

OpenStudy (anonymous):

so 12

OpenStudy (anonymous):

Correct. The smallest number that is divisible by both 4 and 6 is 12, That means the number part we need for the LCD is 12

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

Now let's look at the x's. There is an x^3 and an x. The lowest expression that will divide both x^3 and x is x^3, so we need x^3 for the LCD.

OpenStudy (anonymous):

Now let's look at the y part. You have y and y^2. The simplest expression with y that will divide evenly by y and y^2 is y^2, so you need y^2 in the LCD.

OpenStudy (anonymous):

ok so it is 12x^3y^2

OpenStudy (anonymous):

yes!!

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

:) thank you

OpenStudy (michele_laino):

good job! @aaldia678

OpenStudy (anonymous):

hb the numerators?

OpenStudy (anonymous):

12x^3y^2 is the final answer right. you just needed the LCD

OpenStudy (anonymous):

whatttt

OpenStudy (michele_laino):

we have to simplify the expression above

OpenStudy (anonymous):

btw it is assume no denomiator =0

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

so what is the numerator if we found the denominator

OpenStudy (anonymous):

\[\left(\begin{matrix}2x \\ 0\end{matrix}\right) - \frac{ 3y }{ 0}\] so this is the equation that needs to be simplified?

OpenStudy (anonymous):

sorry, that first part came out weird

OpenStudy (anonymous):

no denominator =0 and thats ok

OpenStudy (michele_laino):

no one of the denominators can be equal to zero @aaldia678

OpenStudy (anonymous):

thx @Michele_Laino

OpenStudy (anonymous):

hmm okay so what would be the full answer?

OpenStudy (michele_laino):

step #1 we have to compute this division: \[\Large \frac{{12{x^3}{y^2}}}{{2x}} = ...?\]

OpenStudy (anonymous):

my question is why is the 12 term the numerator?

OpenStudy (anonymous):

that equals 6x^2y^2

OpenStudy (michele_laino):

ok! now we have to multiply the first numerator by 6x^2y^2, namely, what is: \[\Large 2x \cdot 6{x^2}{y^2} = ...?\]

OpenStudy (anonymous):

12x^3y^2

OpenStudy (michele_laino):

here is wha we are doing: